OpenStudy (anonymous):

A tax exempt charity received a bill for $118.56 which included a 4% sales tax. How much does the charity actually owe?

OpenStudy (anonymous):

118.56*.04=4.74 118.56-4.74=113.82

OpenStudy (anonymous):

OMG! I thought that was it but I'm so nervous I second guessed myself. Thank you so much!

OpenStudy (anonymous):

So did I ...lol! You're welcome! Fan me

OpenStudy (anonymous):

ok trying to figure out how to do that as well...lol

OpenStudy (anonymous):

Click beside screen name

OpenStudy (anonymous):

Become a fan in blue

OpenStudy (anonymous):

This doesn't sound right

OpenStudy (anonymous):

lets say b= original bill and t= tax so b * (.04) = t so [b * (.04) ] + b = 113.82 So we are trying to find b so we factor b from the left and we get b( .04 + 1.00)= 113.82 b( 1.04)= 113.82 b= (113.82)/(1.04) b=(109.4423) we round it to the one cent b= 109.44